An object has weight W on earth. What is the weight of the object if it's on a planet

with 1=6 the mass of the earth and 1=3 the radius of earth?

A. 1

2W

B. W

C. 3

2W

D. 2W



1
Expert's answer
2021-11-21T17:25:36-0500

"W_1=GMm\/r^2"


"W_2=G(M/6)m/(r/3)^2=(GMm/r^2)\cdot(3/2)=3W_1/2"


"C. \ \ 3W/2" . Answer